# Is Limited Information Enough? An Approximate Multi-agent Coverage Control in Non-Convex Discrete Environments

### 08 Jan 2024 v1

@@ Appendix A Proof of Theorem 5 (convergence)
$\begin{array}[]{l}\phi(x^{\prime},\mathcal{P}^{\prime})\\ =\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+[V(x^{\prime},\mathcal{P}^{\prime})-u_{\min}(x^{\prime},\mathcal{P}^{\prime})]_{+}\\ \geq\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})\\ =M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})+\underset{\begin{subarray}{c}i\in\mathcal{N}\setminus\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})\\ \geq\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})\\ =\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})+[V(x,\mathcal{P})-u_{\min}(x,\mathcal{P})]_{+}\\ =\phi(x,\mathcal{P}).\end{array}$ (120)
Note that the equality holds only when $M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})=\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})$. This means that $\phi$ increases as long as the local state in $\mathcal{N}_{i}$ does not satisfy the condition of $Z_{4}$.
@@
$\begin{array}[]{l}M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})\\ =\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+M_{1}(x^{\prime}_{jnearest};\mathcal{P}^{\prime}_{jnearest})\\ \leq\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+V(x^{\prime},\mathcal{P}^{\prime}).\end{array}$ (126)
Since $(x,\mathcal{P})\in Z_{2}\setminus Z_{3}$, $M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})-M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})>u_{\min}(x,\mathcal{P})$. Then,
$\begin{array}[]{l}\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+V(x^{\prime},\mathcal{P}^{\prime})\\ \geq M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})\\ >M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})+u_{\min}(x,\mathcal{P})\\ \geq\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})+u_{\min}(x,\mathcal{P}).\end{array}$ (133)
Then, since $i_{\min}(x,\mathcal{P})\not\in\mathcal{N}_{i}$,
@@
$\begin{array}[]{l}M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})\\ =\underset{\begin{subarray}{c}y_{1},\ldots,y_{k}\in\mathcal{P}_{n_{i}}\end{subarray}}{\max}G(y_{1},\ldots,y_{k};\mathcal{P}_{n_{i}})\\ \geq G(x;\mathcal{P}_{n_{i}}).\\ \geq\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})\\ \end{array}$ (153)
@@
$\begin{array}[]{l}\phi(x^{\prime},\mathcal{P}^{\prime})\\ =\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+[V(x^{\prime},\mathcal{P}^{\prime})-u_{\min}(x^{\prime},\mathcal{P}^{\prime})]_{+}\\ =\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})\\ =M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})+\underset{\begin{subarray}{c}i\in\mathcal{N}\setminus\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})\\ \geq\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})\\ =\phi(x,\mathcal{P}).\end{array}$ (164)
Note that the equality holds only when $M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})=\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})$. This means that $\phi(x,\mathcal{P})$ increases as long as $(x,\mathcal{P})\in Z_{3}$, until $(x,\mathcal{P})$ reaches $Z_{4}$. ∎
@@ Appendix B Proof of Theorem 5 (Approximation ratio)
Since $(x;\mathcal{P})\in Z_{4}\subseteq Z3$, $M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})-M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})\leq u_{\min}(x,\mathcal{P}),\leavevmode\nobreak\ \forall i\in\mathcal{N}$. Also, $M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})=\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})=G(x,\mathcal{P}_{n_{i}}),\leavevmode\nobreak\ \forall i\in\mathcal{N}$. Then, forall $i\in\mathcal{N}$,
$\begin{array}[]{l}M_{1}(x_{n_{i}},\mathcal{P}_{n_{i}})\\ =\underset{\begin{subarray}{c}y\in\mathcal{P}_{n_{i}}\end{subarray}}{\max}G(y,x_{n_{i}};\mathcal{P}_{n_{i}})-G(x_{n_{i}},\mathcal{P}_{n_{i}})\\ =\underset{\begin{subarray}{c}y\in\mathcal{P}_{n_{i}}\end{subarray}}{\max}G(y,x_{n_{i}};\mathcal{P}_{n_{i}})-\underset{\begin{subarray}{c}x_{n_{i}}\in\mathcal{P}_{n_{i}}\end{subarray}}{\max}G(x_{n_{i}},\mathcal{P}_{n_{i}})\\ \leq\underset{\begin{subarray}{c}y_{1},\ldots,y_{n_{i}+1}\in\mathcal{P}_{n_{i}}\end{subarray}}{\max}G(y_{1},\ldots,y_{n_{i}+1};\mathcal{P}_{n_{i}})-\underset{\begin{subarray}{c}y_{1},\ldots,y_{n_{i}}\in\mathcal{P}_{n_{i}}\end{subarray}}{\max}G(y_{1},\ldots,y_{n_{i}};\mathcal{P}_{n_{i}})\\ =M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})-M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})\\ \leq u_{\min}(x,\mathcal{P}).\end{array}$ (176)
Then, $u_{\min}(x,\mathcal{P})\geq\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\max}M_{1}(x_{n_{i}};\mathcal{P}_{n_{i}})=M_{1}(x;\mathcal{P})$. ∎
@@
Because $vor_{i}(x_{S};\mathcal{C})\subseteq vor_{i}(x_{T};\mathcal{C})$. This is because of agents $T-S$. ∎
@@
Appendix C Comparison of algorithms in 1 dimentional setting
We examine the different behavior of algorithms (VVP, SOTA, NBO and the optimal solution, OPT) with simple 1 dimentional setting (Figure 8). Initially, two agents are located at the left end. Agent 0 moves first in all cases. In SOTA, each agent is activated once. The agent first tries to improve the objective function by itself. Then, the agent also tries to communicate with other agents to improve the objective function together. In B, agent 0 tries to move first, but it can’t move because it is blocked by agent 1. Then, agent 0 is deactivated and does not move anymore by itself. Next, agent 1 moves right, stops at its best position, and tries to improve the objective function further by communicating with agent 0. However, the algorithm cannot find any collaborative moves to improve the situation and terminates. Meanwhile, in C, NBO finds an optimal solution because the two agents are neighbors. In this case, the coverage control reaches another optimal solution as in D.
@@ Appendix C Comparison of algorithms in 1 dimentional setting
Figure 8. 1 dimentional setting for comparison of algorithms with $n=2,m=12$. All nodes have the same weight $v_{c}=1$. Agent 0 moves first in all results. A: The initial allocation. B …D: Outcomes of algorithms. $G_{*}$ is the value of the objective function. $G_{OPT}$ is the optimal value. E,F: Outcomes with switching the initial locations of the agents.
Appendix D The Details of Experiments
@@
We evaluate the scalability of the proposed algorithm by changing the size of the space $|C|$ and the number of agents $n$. Also, we set $|\mathcal{C}_{+}|=|C|$ (the top figure in Figure 13). The middle figure shows the runtime until the convergence with different $|C|$. The growing speed of the runtime closely matches the theoretical prediction of $\mathcal{O}(|\mathcal{P}_{n_{i}}|^{2})$, which is polynomial. The bottom figure shows the runtime with different $n$. Surprisingly, the runtime decreases when $n$ increases, because the size of the partitions $|\mathcal{P}_{n_{i}}|$ also decreases. Note that the current implementation uses a single CPU just for theoretical verification, and parallel computation for each agent can reduce the runtime further.

### 08 Apr 2024 v3

@@ Appendix A Proof of Theorem 4 (convergence)
$\begin{array}[]{l}\phi(x^{\prime},\mathcal{P}^{\prime})\\ =\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+[V(x^{\prime},\mathcal{P}^{\prime})-u_{\min}(x^{\prime},\mathcal{P}^{\prime})]_{+}\\ \geq\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})\\ =M_{2}(\emptyset,\mathcal{P}_{S})+\underset{\begin{subarray}{c}i\in\mathcal{N}\setminus S\end{subarray}}{\sum}u_{i}(x,\mathcal{P})\\ \geq\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})\\ =\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})+[V(x,\mathcal{P})-u_{\min}(x,\mathcal{P})]_{+}\\ =\phi(x,\mathcal{P}).\end{array}$ (119)
Note that the equality holds only when $M_{2}(\emptyset,\mathcal{P}_{S})=\underset{\begin{subarray}{c}i\in S\end{subarray}}{\sum}u_{i}(x,\mathcal{P})$. This means that $\phi$ increases as long as the local state in $S$ does not satisfy the condition of $Z_{4}$.
@@
$\begin{array}[]{l}M_{3}(\emptyset,\mathcal{P}_{S})\\ =\underset{\begin{subarray}{c}i\in S\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+M_{1}(x^{\prime}_{jnearest};\mathcal{P}^{\prime}_{jnearest})\\ \leq\underset{\begin{subarray}{c}i\in S\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+V(x^{\prime},\mathcal{P}^{\prime}).\end{array}$ (125)
Since $(x,\mathcal{P})\in Z_{2}\setminus Z_{3}$, $M_{3}(\emptyset,\mathcal{P}_{S})-M_{2}(\emptyset,\mathcal{P}_{S})>u_{\min}(x,\mathcal{P})$. Then,
$\begin{array}[]{l}\underset{\begin{subarray}{c}i\in S\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+V(x^{\prime},\mathcal{P}^{\prime})\\ \geq M_{3}(\emptyset,\mathcal{P}_{S})\\ >M_{2}(\emptyset,\mathcal{P}_{S})+u_{\min}(x,\mathcal{P})\\ \geq\underset{\begin{subarray}{c}i\in S\end{subarray}}{\sum}u_{i}(x,\mathcal{P})+u_{\min}(x,\mathcal{P}).\end{array}$ (132)
Then, since $i_{\min}(x,\mathcal{P})\not\in S$,
@@
$\begin{array}[]{l}M_{2}(\emptyset,\mathcal{P}_{S})\\ =\underset{\begin{subarray}{c}y_{1},\ldots,y_{k}\in\mathcal{P}_{S}\end{subarray}}{\max}G(y_{1},\ldots,y_{k};\mathcal{P}_{S})\\ \geq G(x;\mathcal{P}_{S}).\\ \geq\underset{\begin{subarray}{c}i\in S\end{subarray}}{\sum}u_{i}(x,\mathcal{P})\\ \end{array}$ (152)
@@
$\begin{array}[]{l}\phi(x^{\prime},\mathcal{P}^{\prime})\\ =\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+[V(x^{\prime},\mathcal{P}^{\prime})-u_{\min}(x^{\prime},\mathcal{P}^{\prime})]_{+}\\ =\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})\\ =M_{2}(\emptyset,\mathcal{P}_{S})+\underset{\begin{subarray}{c}i\in\mathcal{N}\setminus S\end{subarray}}{\sum}u_{i}(x,\mathcal{P})\\ \geq\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})\\ =\phi(x,\mathcal{P}).\end{array}$ (163)
Note that the equality holds only when $M_{2}(\emptyset,\mathcal{P}_{S})=\underset{\begin{subarray}{c}i\in S\end{subarray}}{\sum}u_{i}(x,\mathcal{P})$. This means that $\phi(x,\mathcal{P})$ increases as long as $(x,\mathcal{P})\in Z_{3}$, until $(x,\mathcal{P})$ reaches $Z_{4}$. ∎
@@
Note: The proofs above hold even if we replace $S$ with the general definition of neighborhood $\mathcal{N}_{i}$.
@@ Appendix B Proof of Theorem 4 (Approximation ratio)
Since $(x;\mathcal{P})\in Z_{4}\subseteq Z3$, $M_{3}(\emptyset,\mathcal{P}_{S})-M_{2}(\emptyset,\mathcal{P}_{S})\leq u_{\min}(x,\mathcal{P}),\leavevmode\nobreak\ \forall i\in\mathcal{N}$. Also, $M_{2}(\emptyset,\mathcal{P}_{S})=\underset{\begin{subarray}{c}i\in S\end{subarray}}{\sum}u_{i}(x,\mathcal{P})=G(x,\mathcal{P}_{S}),\leavevmode\nobreak\ \forall i\in\mathcal{N}$. Then, forall $i\in\mathcal{N}$,
$\begin{array}[]{l}M_{1}(x_{\{i,j\}},\mathcal{P}_{S})\\ =\underset{\begin{subarray}{c}y\in\mathcal{P}_{S}\end{subarray}}{\max}G(y,x_{\{i,j\}};\mathcal{P}_{S})-G(x_{\{i,j\}},\mathcal{P}_{S})\\ =\underset{\begin{subarray}{c}y\in\mathcal{P}_{S}\end{subarray}}{\max}G(y,x_{\{i,j\}};\mathcal{P}_{S})-\underset{\begin{subarray}{c}x_{\{i,j\}}\in\mathcal{P}_{S}\end{subarray}}{\max}G(x_{\{i,j\}},\mathcal{P}_{S})\\ \leq\underset{\begin{subarray}{c}y_{1},\ldots,y_{3}\in\mathcal{P}_{S}\end{subarray}}{\max}G(y_{1},\ldots,y_{3};\mathcal{P}_{S})-\underset{\begin{subarray}{c}y_{1},y_{2}\in\mathcal{P}_{S}\end{subarray}}{\max}G(y_{1},y_{2};\mathcal{P}_{S})\\ =M_{3}(\emptyset,\mathcal{P}_{S})-M_{2}(\emptyset,\mathcal{P}_{S})\\ \leq u_{\min}(x,\mathcal{P}).\end{array}$ (175)
Then, $u_{\min}(x,\mathcal{P})\geq\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\max}M_{1}(x_{\{i,j\}};\mathcal{P}_{S})=M_{1}(x;\mathcal{P})$. ∎
@@
Because $\mathcal{P}_{i}(x_{S};\mathcal{C})\subseteq\mathcal{P}_{i}(x_{T};\mathcal{C})$. This is because of agents $T-S$. ∎
@@
Note: The proofs above also hold even if we replace $S$ with the general definition of neighborhood $\mathcal{N}_{i}$.
Appendix C Comparison of algorithms in 1 dimensional setting
We examine the different behavior of algorithms (VVP, SOTA, NBO and the optimal solution, OPT) with simple 1-dimensional setting (Figure 9). Initially, two agents are located at the left end. Agent 0 moves first in all cases. In SOTA, each agent is activated once. The agent first tries to improve the objective function by itself. Then, the dagent also tries to communicate with other agents to improve the objective function together. In B, agent 0 tries to move first, but it can’t move because it is blocked by agent 1. Then, agent 0 is deactivated and does not move anymore by itself. Next, agent 1 moves right, stops at its best position, and tries to improve the objective function further by communicating with agent 0. However, the algorithm cannot find any collaborative moves to improve the situation and terminates. Meanwhile, in C, NBO finds an optimal solution because the two agents are neighbors. In this case, the coverage control reaches another optimal solution as in D.
@@ Appendix C Comparison of algorithms in 1 dimensional setting
Figure 9. 1-dimensional setting for comparison of algorithms with $n=2,m=12$. All nodes have the same weight $v_{c}=1$. Agent 0 moves first in all results. A: The initial allocation. B …D: Outcomes of algorithms. $G_{*}$ is the value of the objective function. $G_{OPT}$ is the optimal value. E,F: Outcomes with switching the initial locations of the agents.
Appendix D Details of Experiments
@@
We evaluate the scalability of the proposed algorithm by changing the size of the space $|C|$ and the number of agents $n$. Also, we set $|\mathcal{C}_{+}|=|C|$ (the top figure in Figure 14). The middle figure shows the runtime until the convergence with different $|C|$. The growing speed of the runtime closely matches the theoretical prediction of $\mathcal{O}(|\mathcal{P}_{S}|^{2})$, which is polynomial. The bottom figure shows the runtime with different $n$. Surprisingly, the runtime decreases when $n$ increases, because the size of the partitions $|\mathcal{P}_{S}|$ also decreases. Note that the current implementation uses a single CPU just for theoretical verification, and parallel computation for each agent can reduce the runtime further.